# RoboSpatial: Teaching Spatial Understanding to 2D and 3D Vision-Language Models for Robotics

### v1

@@ Abstract
Spatial understanding is a crucial capability for robots to make grounded decisions based on their environment. This foundational skill enables robots not only to perceive their surroundings but also to reason about and interact meaningfully within the world. In modern robotics, these capabilities are taken on by visual language models, and they face significant challenges when applied to spatial reasoning context due to their training data sources. These sources utilize general-purpose image datasets, and they often lack sophisticated spatial scene understanding capabilities. For example, the datasets do not address reference frame comprehension — spatial relationships require clear contextual understanding, whether from an ego-centric, object-centric, or world-centric perspective, which allow for effective real-world interaction. To address this issue, we introduce RoboSpatial, a large-scale spatial understanding dataset consisting of real indoor and tabletop scenes captured as 3D scans and egocentric images, annotated with rich spatial information relevant to robotics. The dataset includes 1M images, 5K 3D scans, and 3M annotated spatial relationships, with paired 2D egocentric images and 3D scans to make it both 2D and 3D ready. Our experiments show that models trained with RoboSpatial outperform baselines on downstream tasks such as spatial affordance prediction, spatial relationship prediction, and robotics manipulation.
11footnotetext: Corresponding author: song.1855@osu.edu. This work was conducted during the first author’s internship at NVIDIA.
@@ 1 Introduction
Figure 1: An illustration of a model trained on RoboSpatial being used to solve a manipulation task using spatial reasoning. We show that the trained model can enable human-aligned spatial reasoning in the correct reference frame. Furthermore, our chosen spatial relationships enable models to do task grounding, planning, and successful detection.
The introduction of vision-and-language models (VLMs) in modern robotics has helped reconcile general computer vision with robotics control [59]. We have seen these models trained directly on robotics manipulation data [10] which allows robots to directly consume image and task description inputs and act in the real-world. Other works have used generic VLMs to help describe robotics scenes to accomplish specific tasks such as cleaning a table top [16]. We have also seen grasping augmented with text-based description [49]. Moreover, large language models (LLMs) are also important to modern robotics, they have been used extensively to write robot code from user prompts [44, 29] or on high-level planning [1, 45, 21].
Dataset 3D scans Embodied Frames Compatibility Domain #Scans #Images #Spatial QAs
@@
VLMs can recognize objects, classify scenes, and even provide general descriptions that capture high-level attributes. However, despite significant recent advancements, VLMs [40, 34, 30] still fall short in spatial understanding [56, 26, 43]. They struggle with tasks that require interpreting nuanced spatial relationships between objects, such as describing where one object is in relation to another or determining the best location to place an item within a specific condition. For example, while a model might accurately describe a “bowl on the table,” it lacks the ability to reason about where on the table the bowl is, where it should go to ensure accessibility or stability, or how it might fit among other objects. Furthermore, a critical limitation of existing VLM training datasets is their inability to capture reference frame understanding — the way we interpret spatial relationships changes drastically depending on whether we’re viewing from a first-person perspective, focusing on specific objects, or observing the entire scene, all of which are essential for real-world interactions. These limitations underscore the gap between surface-level scene description and the deeper spatial comprehension necessary for intuitive interaction.
@@
To address this gap, various works have proposed methods to teach VLMs about spatial reasoning. Although some recent works have explored training VLMs for spatial understanding, they stop short of meeting the requirements for robotics applications. For example, SpatialVLM [8] and SpatialRGPT [9] train VLMs to answer questions about distances and spatial relationships between objects, advancing spatial understanding at a conceptual level. However, these models are trained on datasets comprised of images from the internet, with annotations generated by perception models. As a result, they struggle to generalize to images captured by robot cameras that often lack identifiable cues for an absolute scale. RoboPoint [57] takes a different approach by training a VLM to produce grounded 2D coordinates, pinpointing object locations or free space within a scene. Yet, due to its reliance on synthetic data, the model lacks an understanding of real-world constraints, such as the amount of space needed to place specific objects or the suitability of surfaces for various items. It is also not able to infer object-centric reference frames, such as identifying what is in front of a car. Most recently, Molmo [12], a suite of vision-language models designed for object centric image-space pointing, shows promise for bridging AI and physical environments. However, our findings reveal limitations in its understanding of real-world constraints, e.g., failing to predict whether the gray bowl can fit in front of the car in Figure 1.
Our hypothesis is that one of the main bottlenecks limiting the effectiveness of VLMs in robotics is the limited training data itself, which is clearly highlighted by Table 1. Therefore, we introduce RoboSpatial, a dataset designed to teach spatial understanding to VLMs with robotic applications in mind. Our approach leverages annotated indoor scene and tabletop RGBD datasets, transforming them into targeted question-answer pairs that probe specific spatial reasoning skills relevant for robotics. We categorize the questions into three types: (1) Spatial configuration, focusing on object-object spatial relationships; (2) Spatial context, addressing object-space spatial relationships; and (3) Spatial compatibility, exploring object-object compatibility. Each question type serves a distinct purpose: (1) The spatial configuration questions focus on binary assessments, asking whether certain spatial relationships exist between pairs of objects, such as “Is the mug to the left of the laptop?” (2) Spatial context questions are designed as point-prediction tasks, challenging the model to determine appropriate locations within free space where an object can be placed, such as “Where on the table can I put the plate?” Finally, (3) the spatial compatibility questions also use a binary format to address compatibility between objects, such as “Can the chair be placed in front of the table?”
@@
To enhance the model’s ability to interpret spatial instructions from different perspectives, each question-answer pair in RoboSpatial is posed from three distinct reference perspectives/frames: (a) Ego-centric from the observer’s perspective at the camera pose, (b) Object-centric based on a reference frame attached to the focal object, and (c) World-centric grounded in a global world frame. This multi-frame approach enables models to handle complex spatial instructions more flexibly, preparing them to better generalize to dynamic robotic contexts. Applying our methodology to existing indoor scene and tabletop datasets, we generate both a comprehensive training dataset and a benchmark for spatial question answering in robotics. RoboSpatial contains around 1M images, 5k 3D scans, and 3M annotated spatial relationships, with paired 2D egocentric images and 3D scans to make it both 2D and 3D ready, see Figure 1 for an example taken from our dataset.
To validate the effectiveness of our dataset, we conduct comprehensive experiments with multiple state-of-the-art (SOTA) 2D and 3D VLMs. Results show that training VLMs on RoboSpatial significantly enhances their spatial reasoning capabilities, with models trained on our dataset outperforming existing models on our test split. To further demonstrate the robustness of RoboSpatial, we assess the trained VLMs on three additional downstream robotic tasks: RoboSpatial-Home, , a high-quality manually annotated indoor scene question-answering dataset, BLINK-Spatial [17], and real-world robot manipulation. These tasks directly evaluate the models’ ability to apply spatial understanding in practical scenarios, such as rearranging objects or responding to context-specific spatial questions within an indoor environment. Additionally, we test our models on an external benchmark, BLINK-Spatial [17], to measure generalization to unseen spatial reasoning tasks beyond our dataset. Leveraging RoboSpatial’s 3D-ready design, we also compare the performance of 2D and 3D VLMs directly. These results emphasize the strengths of RoboSpatial in advancing spatial understanding and reinforce the benefits of 3D-based training for enhancing the spatial reasoning capabilities of VLMs in robotic applications.
@@
  • A new training dataset and benchmark, RoboSpatial and RoboSpatial-Home, comprising images and 3D scans paired with spatial questions and answers, designed with unique features including multiple reference frames, object-object spatial relationships, object-space relationships, and object compatibility. We plan to publicly release our data and code for generating the dataset from 3D annotated scenes.
  • Training of VLMs on RoboSpatial, resulting in enhanced spatial reasoning capabilities that significantly outperform prior VLMs on natural language-specified robot manipulation tasks and indoor scene question answering.
@@ 2 Related Work
VLMs for Robotics. Vision-language models (VLMs) have emerged as pivotal tools in robotics, enabling systems to interpret and act upon complex visual and textual information. By integrating visual perception with language understanding, VLMs facilitate more intuitive human-robot interactions and enhance autonomous decision-making capabilities. Recent advancements have demonstrated the potential of VLMs in various robotic applications. For instance, vision-and-language action models (VLAs) [27, 4, 39] enable robots to interpret and execute complex instructions and output executable robot actions. Additionally, VLMs like GPT-4v [40] have been utilized for high-level task planning [51], allowing robots to generate detailed action sequences from natural language instructions. Furthermore, VLMs have been used for keypoint prediction [22, 16, 38], error analysis [14], grasp pose prediction [19]. Despite these advancements, integrating VLMs [5, 9, 57] into robotic systems presents challenges. One significant hurdle is the need for precise spatial reasoning to navigate and manipulate objects effectively. While VLMs excel in understanding and generating language, their ability to comprehend and reason about spatial relationships in dynamic environments remains limited [56, 55, 53]. Therefore, RoboSpatial aims to tackle this gap by presenting a large scale pretraining and evaluation setup for teaching spatial understanding to VLM for robotics.
Figure 2: Overview of the RoboSpatial dataset. We automatically generate spatial relationship annotations from existing datasets with 3D point clouds, egocentric images, and 3D bounding box annotations. We create question/answer pairs covering three classes of spatial relationships, three spatial reference frames, and both binary (yes/no) and numeric (e.g. 2D image points) answers. From 1M images and 1.4K scans, we generate over 2M spatial question/answer pairs.
Spatial Understanding with VLMs. Spatial understanding has been implicitly and explicitly part of various vision and question answering tasks [17, 3, 24, 46, 2, 28, 25, 23]. While many benchmarks and methods have been proposed, they often come with limitations: some focus exclusively on simulations [48] or generic images [33, 41, 9, 8, 17, 26, 43], others are difficult to evaluate [48, 13], rely on complete 3D scans [58, 37, 36, 32], do not consider reference frames [58, 37, 36, 32, 8, 9, 17], or use data generated by LLMs that may lack sufficient accuracy [32]. Furthermore, they often fail to address actionable, robotics-relevant spatial relationships such as spatial compatibility and context [13, 17, 54, 43, 26].
@@
To achieve this, we have developed and are planning to open-source a large-scale 2D/3D ready pretraining dataset, an automated data annotation pipeline, and trained models. We further show how our dataset can be used to teach spatial reasoning to a suite of vision-language models (VLMs) in in-domain and out-of-domain spatial reasoning datasets. We hope these resources lower the barrier to entry for exploring spatial understanding tailored to robotics.
@@ 3 Approach
In this section, we begin by explaining our selection of three spatial relationships: spatial configuration, spatial context, and spatial compatibility. Next, we provide an overview of the data generation pipeline used to construct RoboSpatial. An overview of the RoboSpatial is provided in Figure 2.
@@ 3.1 Spatial Relationships
Our dataset is organized around three core spatial relationships that we believe to address the essential aspects of spatial reasoning for robotic tasks: object configuration, object context, and object compatibility. Configuration enables robots to understand and interpret the relative positioning of objects, which is crucial for directing navigation, manipulation, and interaction within complex environments. Context allows robots to assess the relationship between objects and their surrounding space, facilitating the identification of empty or occupied areas and aiding in efficient path planning and obstacle avoidance. Compatibility focuses on whether objects can coexist or interact without conflict in a given space, which is vital for object placement, assembly, and operational safety. Together, these spatial relationships provide a more nuanced and practical framework for robotic applications than metrics like distance—which is hard to normalize across different scales, environments, and tasks—thereby enabling robots to perform complex tasks with greater reliability.
@@ 3.2 Dataset Generation
The goal of the data construction pipeline is to generate a high-accuracy spatial relationship dataset with minimal human intervention by use of carefully constructed heuristics. Our data generation process takes as input a source scene dataset $\mathcal{D}_{s}$ that contains RGB views, camera poses, and object annotations in form of text labels and oriented 3D bounding boxes. The output is a dataset $\mathcal{D}$ where each datum $d_{i}=\langle I_{i},q_{i},a_{i},l_{i}\rangle$ consists of an image $I_{i}$, a question $q_{i}$, an answer $a_{i}$, and a reference frame label $l_{i}\in\{\textit{ego},\textit{world},\textit{object}\}$. Each question belongs to one of the three core spatial reasoning categories or a fourth category for object grounding (Table 2).
Our data generation process consists of two stages, (1) Spatial Relation Extraction, and (2) Question-Answer Generation.
3.2.1 Spatial Relation Extraction
@@
In the spatial relation extraction stage, we analyze the dataset to discover a set of spatial relations between objects or points in free space. Each spatial relation can be described as $\langle I_{i},a_{i},t_{i},r_{i},l_{i}\rangle$, where $I_{i}$ is an image, $a_{i}$ is an anchor object, $t_{i}$ is a target object or a target free-space point, $s_{i}$ is size measure based on an object to be manipulated, $r_{i}\in\{\textit{left},\textit{right},\textit{above},\textit{below},\textit{front},\textit{behind}\}$ is a spatial relation preposition, and $l_{i}\in\{\textit{ego},\textit{world},\textit{object}\}$ is a reference frame label.
We use oriented 3D bounding box annotations to automatically generate spatial relationships. Since the 3D bounding boxes contain information about orientation (which direction the object is facing) and location in the world (including elevation), we then utilize a heuristic to automatically extract spatial relationships between objects and space.
Spatial Configuration: Spatial configuration relationships involve determining whether one object is in a certain spatial direction relative to an anchor object. We first choose an anchor object and generate spatial relationships with all the objects that appear in the image. For each of the three reference frames, we select the direction indicated by the spatial relation. For example, if the spatial relation is (in front, object frame), we only select objects that are in the forward direction from the oriented bounding box. To avoid ambiguity, we only consider objects that appear once in the current frame and generate relationships between all object pairs in the image. Answers for this relationship are binary (True/False).
Spatial Context: Spatial context relationship refers to a task where the model is required to determine points (x,y coordinates in the camera frame, also known as image space) relative to the spatial relationship with the anchor object. One example of such a relationship is identifying points that are in front of another object. Specifically, we construct a top-down map of the environment using the annotated 3D bounding boxes and randomly sample points in empty areas that are a set distance from the object. This distance could be based on the size of another object that is to be manipulated and placed in relation to the object. We apply the same reference frame logic as described for the spatial configuration heuristic. Answers to this relationship are numeric, i.e., a list of 2D coordinates that fit the criteria.
Spatial Compatibility: Spatial compatibility is an extension of the spatial context relationship where the model needs to reason about whether a referenced object can fit in a region indicated relative to the anchor object. We use the same top-down map constructed for spatial context generation and simulate placing an object in the map using its bounding box size. We apply the same reference frame logic to define the area we should simulate placing an object. Answers for this relationship are binary (True/False).
Question Types Category Answer type
Spatial Relations Spatial Configuration True/False
Spatial Context 2D points
Spatial Compatibility True/False
Grounding Object Referring 2D bounding box
Table 2: Question Taxonomy.

### v2

@@ Abstract
Spatial understanding is a crucial capability that enables robots to perceive their surroundings, reason about their environment, and interact with it meaningfully. In modern robotics, these capabilities are increasingly provided by vision-language models. However, these models face significant challenges in spatial reasoning tasks, as their training data are based on general-purpose image datasets that often lack sophisticated spatial understanding. For example, datasets frequently do not capture reference frame comprehension, yet effective spatial reasoning requires understanding whether to reason from ego-, world-, or object-centric perspectives. To address this issue, we introduce RoboSpatial, a large-scale dataset for spatial understanding in robotics. It consists of real indoor and tabletop scenes, captured as 3D scans and egocentric images, and annotated with rich spatial information relevant to robotics. The dataset includes 1M images, 5k 3D scans, and 3M annotated spatial relationships, and the pairing of 2D egocentric images with 3D scans makes it both 2D- and 3D- ready. Our experiments show that models trained with RoboSpatial outperform baselines on downstream tasks such as spatial affordance prediction, spatial relationship prediction, and robotics manipulation.
11footnotetext: Corresponding author: song.1855@osu.edu. This work was partly done during the first author’s internship at NVIDIA.
@@ 1 Introduction
Figure 1: RoboSpatial dataset facilitates 3D spatial reasoning for robotic manipulation. This illustration demonstrates how a model trained on RoboSpatial enables human-aligned spatial reasoning within the correct reference frame, supporting task grounding, planning, and detection for manipulation tasks.
The rise of vision-language models (VLMs) has opened new opportunities for agents to interpret and act upon the visual world using natural language. VLMs have been adopted across a range of embodied settings, notably in robotics and augmented reality (AR). In robotics, they have enabled grounded scene understanding [13, 66], manipulation [7], and policy code generation [48, 29], while in AR, they support tasks like object labeling [50], action recognition [17, 16], and temporal grounding [22].
Dataset 3D scans Embodied Ref. frames Compatibility Domain #Scans #Images #Spatial QAs
@@
VLMs can recognize objects, classify scenes, and even provide general descriptions that capture high-level attributes. However, despite significant recent advancements, VLMs [42, 35, 30] still fall short in spatial understanding [61, 26, 47, 38]. They struggle with tasks that require interpreting nuanced spatial relationships between objects, such as describing where one object is in relation to another or determining the best location to place an item within a specific condition. For example, while a model might accurately describe a “bowl on the table,” it lacks the ability to reason about where on the table the bowl is, where it should go to ensure accessibility or stability, or how it might fit among other objects. Furthermore, a critical limitation of existing VLM training datasets is their inability to capture reference frame understanding (ref. frame) — the way we interpret spatial relationships changes drastically depending on whether we’re viewing from a first-person perspective, focusing on specific objects, or observing the entire scene, all of which are essential for real-world interactions.
@@
These limitations highlight an ongoing challenge: bridging the gap between surface-level scene description and the deeper spatial comprehension necessary for intuitive interaction. Several recent efforts aim to address this by explicitly training VLMs on spatial reasoning tasks, yet many fall short of the demands posed by embodied or robotics settings. For example, SpatialVLM [5] and SpatialRGPT [6] train VLMs to answer questions about distances and spatial relationships between objects, advancing spatial understanding at a conceptual level. However, these models are trained on datasets comprised of images from the internet, with annotations generated by perception models. As a result, they struggle to generalize to embodied images—those captured by robot cameras within real-world environments, which often lack identifiable cues for an absolute scale. Pointing models, such as RoboPoint [62] and more recently Molmo [9], take a different approach by training VLMs to produce grounded 2D coordinates that pinpoint object locations or free space within a scene. However, these models lack understanding of real-world constraints, such as inferring object-centric reference frames for perspective-invariant reasoning, or accounting for the space required to place various objects. This results, for example, in failing to predict whether the gray bowl can fit in front of the car in Figure 1.
This paper hypothesizes that a primary bottleneck limiting the effectiveness of VLMs in robotics is the scarcity of suitable training data, as highlighted by Table 1. To address this, we introduce RoboSpatial, a dataset designed specifically to facilitate spatial understanding in VLMs for robotic applications. The proposed approach leverages annotated indoor scene and tabletop RGBD datasets, transforming them into targeted question-answer pairs designed to probe spatial reasoning skills critical for robotics.
We categorize the questions into three types, each serving a distinct purpose. Spatial context focuses on identifying empty space or support surfaces in the environment that can accommodate other objects. These questions are formulated as point-prediction tasks, challenging the model to determine appropriate locations within free space where an object can be placed—for example, “Where on the table can I put the plate?” Spatial compatibility builds on the identified empty space to assess whether a given area can feasibly support the placement of a specific object, ensuring sufficient size and fit. These questions are posed in a binary format, such as “Can the chair be placed in front of the table?” Spatial configuration examines the relative spatial relationships between two objects. These questions use a binary format to determine whether a spatial relation holds, such as “Is the mug to the left of the laptop?”
@@
To enhance the model’s ability to interpret spatial instructions from different perspectives, each question-answer pair in RoboSpatial is posed from three distinct reference perspectives/frames: (a) Ego-centric from the observer’s perspective at the camera pose, (b) World-centric grounded in a global world frame, and (c) Object-centric based on a reference frame attached to the focal object. This multi-frame approach enables models to handle complex spatial instructions more flexibly, preparing them to better generalize to dynamic robotic contexts. Applying our methodology to existing indoor scene and tabletop datasets, we generate both a comprehensive training dataset and a benchmark for spatial question answering in robotics. RoboSpatial contains around 1M images, 5k 3D scans, and 3M annotated spatial relationships, with paired 2D egocentric images and 3D scans to make it both 2D- and 3D- ready.
To validate the effectiveness of RoboSpatial, comprehensive experiments were conducted using multiple state-of-the-art (SOTA) 2D and 3D VLMs. Results demonstrate that models trained on RoboSpatial exhibit significantly improved spatial reasoning capabilities, consistently outperforming baseline methods on the evaluation benchmark RoboSpatial-Val, a held-out validation subset derived from the heuristically generated RoboSpatial dataset. To further assess the generalization and robustness of these trained VLMs, additional evaluations were conducted using three complementary benchmarks: RoboSpatial-Home, a manually collected dataset consisting of paired RGB and depth images, and two external benchmarks, BLINK-Spatial [15] and SpatialBot [2]. These benchmarks rigorously test spatial reasoning skills in practical robotic tasks, including object rearrangement and contextual question answering in indoor environments, while also examining the models’ capacity to generalize to novel spatial reasoning scenarios beyond the original training data. Across all benchmarks, models trained on RoboSpatial consistently outperformed baseline methods, demonstrating the broad utility of the dataset. Leveraging the 3D-ready design of RoboSpatial, direct comparisons between the spatial reasoning performance of 2D and 3D VLMs were also performed. Although initial results indicate potential advantages for 3D models, differences in pretraining data and base LLM architectures among models render the comparison inconclusive. RoboSpatial is specifically designed to support both 2D and 3D research directions, enabling future studies to address these differences more conclusively.
@@
  • A new training dataset, RoboSpatial, comprising images and 3D scans paired with spatial questions and answers, accompanied by an evaluation benchmark, RoboSpatial-Val, a held-out validation set. Additionally, we introduce RoboSpatial-Home, a manually collected and annotated dataset designed specifically for assessing real-world spatial reasoning in indoor environments. These datasets uniquely incorporate multiple reference frames, object-object spatial relationships, object-space relationships, and object compatibility. We make the data and code for generating the dataset from 3D annotated scenes publicly available [^1] https://chanh.ee/RoboSpatial/.
  • VLMs trained on RoboSpatial demonstrate superior spatial reasoning, outperforming SOTA baselines on language-guided robot manipulation and indoor scene question answering.
@@ 2 Related Work
VLMs for Robotics. Vision-language models (VLMs) have emerged as pivotal tools in robotics, enabling systems to interpret and act upon complex visual and textual information. By integrating visual perception with language understanding, VLMs facilitate more intuitive human-robot interactions and enhance autonomous decision-making capabilities. Recent advancements have demonstrated the potential of VLMs in various robotic applications. For instance, vision-language-action models (VLAs) [27, 65, 41] enable robots to interpret and execute complex instructions and output executable robot actions. Additionally, VLMs like GPT-4v [42] have been utilized for high-level task planning [55], allowing robots to generate detailed action sequences from natural language instructions. Furthermore, VLMs have been used for keypoint/mask prediction [21, 59, 40], error analysis [11, 49], grasp pose prediction [19]. Despite these advancements, integrating VLMs [2, 6, 62] into robotic systems presents challenges. One significant hurdle is the need for precise spatial reasoning to navigate and manipulate objects effectively. While VLMs excel in understanding and generating language, their ability to comprehend and reason about spatial relationships in dynamic environments remains limited [61, 60, 57]. Therefore, RoboSpatial aims to tackle this gap by presenting a large scale pretraining and evaluation setup for teaching spatial understanding to VLM for robotics.
Figure 2: Overview of the RoboSpatial dataset. We automatically generate spatial relationship annotations from existing datasets with 3D point clouds, egocentric images, and 3D bounding box annotations. We create question/answer pairs covering three classes of spatial relationships, three spatial reference frames, and both binary (yes/no) and numeric (e.g. 2D image points) answers. From 1M images and 5k scans, we generate over 3M spatial question/answer pairs.
Spatial Understanding with VLMs. Spatial understanding has been implicitly and explicitly part of various vision and question answering tasks [15, 1, 24, 51, 46, 28, 25, 23]. While many benchmarks and methods have been proposed, they often come with limitations: some focus exclusively on simulations [53] or generic images [33, 43, 6, 5, 15, 26, 47, 44], others are difficult to evaluate due to their reliance on free-form text outputs [53, 10, 32], some rely on complete 3D scans [64, 39, 37, 32], and others do not account for reference frames [64, 39, 37, 32, 5, 6, 15, 44]. Furthermore, many fail to address actionable, robotics-relevant spatial relationships such as spatial compatibility and context [10, 15, 58, 47, 26, 32, 44].
@@
To achieve this, we have developed a large-scale 2D/3D ready training dataset using our automated data generation pipeline. We further show how RoboSpatial can be used to teach spatial reasoning to a suite of vision-language models (VLMs) in in-domain and out-of-domain spatial reasoning datasets. We hope these resources lower the barrier to entry for exploring spatial understanding tailored to robotics.
@@ 3 Approach
We begin by explaining the selection of three spatial relationships: spatial context, spatial compatibility, and spatial configuration. Next, we describe the data generation pipeline used to construct the RoboSpatial. Figure 2 provides an overview of the dataset.
@@ 3.1 Spatial Relationships
The dataset is organized around three core spatial relationships that we believe address the essential aspects of spatial reasoning for robotic tasks: spatial context, spatial compatibility, and spatial configuration. Context allows robots to assess the relationship between objects and their surrounding space, facilitating the identification of empty or occupied areas, which is relevant for downstream applications such as path planning and obstacle avoidance. Compatibility focuses on whether objects can coexist or interact without conflict in a given space, which is vital for object placement, assembly, and operational safety. Configuration enables robots to understand and interpret the relative positioning of objects, which is crucial for directing navigation, manipulation, and interaction within complex environments. Together, these spatial relationships provide a more nuanced and practical framework for robotic applications than metrics like distance—which is hard to normalize across different scales, environments, and tasks—thereby enabling robots to perform complex tasks with greater reliability.
@@ 3.2 Dataset Generation
The goal of the data construction pipeline is to generate a large-scale, high-accuracy spatial relationship dataset with minimal human intervention, using automatic heuristics grounded in 3D geometry and 2D image views.
@@
The pipeline takes as input a scene dataset $\mathcal{D}_{s}$ that contains RGB images, camera poses (both extrinsic and intrinsic parameters), and oriented 3D bounding box annotations with semantic object labels. The output is a spatial reasoning dataset $\mathcal{D}$, where each entry $d_{i}=\langle I_{i},q_{i},a_{i},l_{i}\rangle$ consists of an image $I_{i}$, a question $q_{i}$, an answer $a_{i}$, and a reference frame label $l_{i}\in\{\textit{ego},\textit{world},\textit{object}\}$. Each question is derived from one of three spatial reasoning categories: spatial configuration, spatial context, or spatial compatibility. To support reliable object reference resolution, we also generate an auxiliary object grounding dataset that links object descriptions to 2D bounding boxes.
To improve clarity and reproducibility, we describe the data generation pipeline in two main stages. We also separate the reasoning logic used for extracting 3D relationships from that used for generating 2D image-space targets.
3.2.1 Stage 1: 3D Spatial Relation Extraction
@@
The first stage involves extracting spatial relationships between objects or between objects and free space, based on 3D geometry. Each spatial relation is defined as $s_{i}=\langle I_{i},a_{i},t_{i},r_{i},l_{i}\rangle$, where $I_{i}$ is the source image, $a_{i}$ is the anchor object, $t_{i}$ is the target object or a sampled point in free space, $r_{i}\in\{\textit{left},\textit{right},\textit{above},\textit{below},\textit{front},\textit{behind}\}$ is the relation preposition, and $l_{i}\in\{\textit{ego},\textit{world},\textit{object}\}$ denotes the reference frame.
We use oriented 3D bounding boxes, provided by the source dataset, to compute spatial relationships. Each bounding box includes both the 3D location and heading of the object. The object’s orientation is defined by the heading vector of the bounding box, aligned with the object’s front-facing direction. Using this orientation, we determine the appropriate directional region (e.g., front, left) relative to the reference frame. For instance, a relation such as “in front of (anchor object) (object frame)” refers to the positive direction along the anchor object’s heading vector. These relationships are calculated independently for each of the three reference frames: the world frame is aligned with the dataset-level coordinate system; the ego frame is defined by the camera pose (i.e., camera-centered); and the object frame is defined by the local orientation of the anchor object.
The camera extrinsics are used to transform coordinates between reference frames. Although the method does not require point clouds or meshes, it relies on camera intrinsics and extrinsics to project between 2D and 3D and to ensure consistent reference frame reasoning. For each spatial configuration task, we evaluate all visible object pairs that appear uniquely in the image, avoiding duplicate instances to minimize ambiguity. The resulting relationships are binary (True/False) and specify whether the spatial condition holds for the given object pair.
3.2.2 Stage 2: 2D Spatial Point and Region Sampling
In the second stage, we generate 2D image-space annotations for spatial context and spatial compatibility tasks. These rely on the 3D bounding box layout and calibrated camera parameters to map spatial relationships into image coordinates.
For spatial context, we construct a top-down occupancy map of the scene by marking regions occupied by 3D bounding boxes. We then randomly sample 3D points in empty space that lie in a specified directional relation to the anchor object, following the same frame-dependent heuristics as in the configuration task. These points are projected into the image plane using the camera intrinsics. To ensure the points are valid, we filter out samples that are obstructed or occluded based on line-of-sight from the camera. Specifically, we perform raycasting from the camera center to each sampled 3D point, and discard points whose rays intersect any occupied bounding box volumes before reaching the target location. The final answer is a list of 2D $(x,y)$ image coordinates that satisfy the spatial context constraint.
Spatial compatibility extends this idea by checking whether a target object can fit within the sampled region. We simulate placing a virtual bounding box, matching the size of the target object, at the candidate location on the ground plane. A region is considered compatible if the simulated placement does not intersect with any existing bounding boxes in the scene and provides at least a 10 cm margin along each axis. The simulation allows for translation and in-plane rotation of the object. The answer to this task is binary (True/False), indicating whether the region can accommodate the object.